OpenStudy (anonymous):

What is 145/7 in proper fraction form? PLease help!

OpenStudy (anonymous):

\[20\frac{ 5 }{ 7 }\]

OpenStudy (anonymous):

\[\frac{ 20*7+5 }{ 7 }=\frac{ 140+5 }{ 7 }=\frac{ 145 }{ 7 }\]

OpenStudy (cwrw238):

divide 145 / 7 = 20 remainder 5 and write as lonliness has

OpenStudy (anonymous):

Okay thank you!

OpenStudy (cwrw238):

yw